OpenStudy (anonymous):

Two electric motors drive two elevators of equal mass in a three-story building 10 meters tall. Each elevator has a mass of 1,000 kg. The first elevator can do this work in 5.0 seconds. Calculate the power output of the first motor.

OpenStudy (anonymous):

P ouput = W/t = m g h /t where m is mass of two electric motor, g is accelertion of gravity (10 m/s^2), h is building tall, and t is time. So P output = (1000 kg x 10 x 10)/5 Poutput = 20 000 Watt